# Explaining Time Series Classifiers with PHAR: Rule Extraction and Fusion from Post-hoc Attributions

### v3

@@ 5 Conclusion
Our framework consolidates knowledge from multiple explainers (RQ3). Unlike bagging or boosting, our post-hoc fusion refines single or multiple explainers into a coherent rule set, aligning with expert system principles. Fusion improves rule reliability, reduces ambiguity, and enhances decision-making. Moreover, PHAR supports sparsity optimization. We measure sparsity with the rule-length metric $F(n)$ and report both its mean and median. This is important given human working-memory limits (the 7 $\pm$ 2 principle) [19], as discussed earlier when defining the objective function. Our method enables this, when desired, by offering resolution modes that promote concise, non-overlapping rules, and by using fusion methods such as lasso and intersection. This is also illustrated by the ECG200 example discussed in our qualitative evaluation.
@@ 5.2 Future work
    One of the important directions is extending PHAR to counterfactual explanation generation, by identifying and presenting the nearest time series belonging to an alternative class. This would allow Domain Experts to specify transitions from the current predicted class to any arbitrary target class (e.g., in predictive maintenance or clinical decision support), thereby providing actionable "WHAT-IF" insights.
@@
    Integrating expert feedback could refine rule usability, bridging automated rule extraction with human interpretability. One approach is incorporating symbolic reasoning systems like HeaRTDroid [8]  [^9] https://heartdroid.re/, a lightweight rule engine for expert systems, enabling a hybrid approach that merges statistical learning with structured decision-making.
@@
    Usage of advanced visualization methods, such as interactive overlays or multi-dimensional interval representations could be beneficial. Related direction is DeepVATS [26]  [^10] https://github.com/vrodriguezf/deepvats, which integrates DL with interactive visualization of latent space projections for TS, similar to TensorFlow Embeddings Projector, aiding in anomaly detection and structure analysis.
@@
Acknowledgements This paper is part of a project that has received funding from the European Union’s Horizon Europe Research and Innovation Programme under Grant Agreement No. 101120406. The paper reflects only the authors’ view and the European Commission is not responsible for any use that may be made of the information contained herein.
Contribution of Maciej Mozolewski for the research for this publication has been supported by a grant from the Priority Research Area (DigiWorld) under the Mark Kac Center for Complex Systems Research Strategic Programme Excellence Initiative at Jagiellonian University.
We thank all contributors of open-source libraries (e.g., TensorFlow, Optuna) and the maintainers of the UCR/UEA Time Series Classification Archive for curating and providing the datasets. We used OpenAI’s GPT to assist in writing and refining the text and generating code snippets, while all conceptual ideas and research contributions were developed by the authors.
@@ B Supplementary Materials
Friedman and Nemenyi: To compare multiple methods statistically, we applied the Friedman test, followed by post-hoc Nemenyi comparisons. The Friedman test results provide insights into the statistical significance of differences across methods. Table 22 reports results at the method level, considering different fusion strategies (e.g., Baseline, Best, Intersection, Weighted). Table 23 presents results for all rule sets, accounting for the combinations of explainers they comprise, both single and compound. For pairwise comparisons, we employed the Nemenyi test. The first set of tables, from Table 24 to Table 31, evaluates fusion combinations in terms of explainer types. The second set, from Table 32 to Table 39, focuses on comparisons across fusion methods.

### v4

@@ 5 Conclusion
Our framework consolidates knowledge from multiple explainers (RQ3). Unlike bagging or boosting, our post-hoc fusion refines single or multiple explainers into a coherent rule set, aligning with expert system principles. Fusion improves rule reliability, reduces ambiguity, and enhances decision-making. Moreover, PHAR supports sparsity optimization. We measure sparsity with the rule-length metric $F(n)$ and report both its mean and median. This is important given human working-memory limits (the 7 $\pm$ 2 principle) [Miller1956], as discussed earlier when defining the objective function. Our method enables this, when desired, by offering resolution modes that promote concise, non-overlapping rules, and by using fusion methods such as lasso and intersection. This is also illustrated by the ECG200 example discussed in our qualitative evaluation.
@@ 5.2 Future work
    One of the important directions is extending PHAR to counterfactual explanation generation, by identifying and presenting the nearest time series belonging to an alternative class. This would allow Domain Experts to specify transitions from the current predicted class to any arbitrary target class (e.g. in predictive maintenance or clinical decision support), thereby providing actionable "WHAT-IF" insights.
@@
    Integrating expert feedback could refine rule usability, bridging automated rule extraction with human interpretability. One approach is incorporating symbolic reasoning systems like HeaRTDroid [heartdroid_2019]  [^9] https://heartdroid.re/, a lightweight rule engine for expert systems, enabling a hybrid approach that merges statistical learning with structured decision-making.
@@
    Usage of advanced visualization methods, such as interactive overlays or multi-dimensional interval representations could be beneficial. Related direction is DeepVATS [RODRIGUEZFERNANDEZ2023110793]  [^10] https://github.com/vrodriguezf/deepvats, which integrates DL with interactive visualization of latent space projections for TS, similar to TensorFlow Embeddings Projector, aiding in anomaly detection and structure analysis.
@@
Acknowledgements This article is part of a project that has received funding from the European Union’s Horizon Europe Research and Innovation Programme under Grant Agreement No. 101120406. The article reflects only the authors’ view and the European Commission is not responsible for any use that may be made of the information contained herein.
The contribution of Maciej Mozolewski for the research for this publication has been supported by a grant from the Priority Research Area (DigiWorld) under the Mark Kac Center for Complex Systems Research Strategic Programme Excellence Initiative at Jagiellonian University.
We thank all contributors of open-source libraries (e.g. TensorFlow, Optuna) and the maintainers of the UCR/UEA Time Series Classification Archive for curating and providing the datasets. We used OpenAI’s GPT to assist in writing and refining the text and generating code snippets, while all conceptual ideas and research contributions were developed by the authors.
[…]
Declaration of Generative AI and AI-Assisted Technologies in the Manuscript Preparation Process During the preparation of this manuscript, the authors used ChatGPT to refine the language, improve the readability, and check grammatical consistency, and to draft selected code snippets, typing, and docstrings. All such suggestions and code were reviewed, tested, and, when needed, rewritten by the authors, who assume full responsibility for the final text and software. ChatGPT did not generate any numerical results, metrics, tables, or figures; all results were obtained by running our own code on the described data and models.
@@ B Supplementary Materials
Friedman and Nemenyi: To compare multiple methods statistically, we applied the Friedman test, followed by post-hoc Nemenyi comparisons. The Friedman test results provide insights into the statistical significance of differences across methods. Table 22 reports results at the method level, considering different fusion strategies (e.g. Baseline, Best, Intersection, Weighted). Table 23 presents results for all rule sets, accounting for the combinations of explainers they comprise, both single and compound. For pairwise comparisons, we employed the Nemenyi test. The first set of tables, from Table 24 to Table 31, evaluates fusion combinations in terms of explainer types. The second set, from Table 32 to Table 39, focuses on comparisons across fusion methods.